There are 34 learners in a grade 9 class. The class register shows that there are 4 more girls than boys. How many boys are there in the class?
Found 3 solutions by jim_thompson5910, adamhen894, ikleyn:
Answer by jim_thompson5910(35256)


Answer: 15 boys in the class
Extra info: there are 19 girls (four more than the boys). Note how 15+19 = 34
-----------------------------------------------------------------------------
-----------------------------------------------------------------------------

Explanation:

Let
x = number of boys
y = number of girls

There are 34 students meaning that
x+y = 34

There are 4 more girls than boys, so
y = x+4

Plug the second equation into the first and solve for x
x+y = 34
x+x+4 = 34
2x+4 = 34
2x+4-4 = 34-4
2x = 30
2x/2 = 30/2
x = 15
There are 15 boys.

We have the answer so we can stop here. As extra bit of info, you can plug x = 15 into the equation y = x+4 to find the number of girls
y = x+4
y = 15+4
y = 19
There are 19 girls.

Answer by adamhen894(15)

first one is to find what question is asking for.
x: number of boys in class
boys : x
girls:x+4 (four more girls than boys)
boys + girls = total students
set up the equation
x + (x+4) = 34
x = 15 (number of boys )
x+4 = 19 ( number of girls)
if you have better way to solve it, please tell me, i would love to hear!!
adamchen894@gmail.com

Answer by ikleyn(52786)

.
Ask 4 girls to stay away for 1 minute.

Then you will have 30 students and equal number of boys and girls.

Hence . . .

Should I continue further? Or everything is just clear to you ?

Now allow these 4 girls to return to the class . . .